If P+1/p=8, then find the value of (p-8)^11+(1/p^11) =?


1
Expert's answer
2021-01-27T03:54:23-0500

Given P+1/p=8 eq( 1)

P-8=-1/p

(P-8)11+(1/p11)

=(-1/p)11+(1/p11)

0 Ans